OpenStudy (anonymous):

sin^6(x) integration

OpenStudy (anonymous):

sin^6(x) = (sin^2(x))^3 sin^2(x) = (1 - cos(2x))/2 expand the polynomial. after expanding: remember that cos^2(x) = (1 + cos(2x))/2 and that to integrate sin^3(x): you'll look at it as sin(x)*(1 - cos^2(x)) , taking u = cos(x)